OpenStudy (anonymous):

A cylinder has a volume of 175pi cubic units and a height of 7 units. The diameter of the cylinder is 5 units 10 units 12.5 units

OpenStudy (anonymous):

@dtan5457

OpenStudy (danjs):

175 = pi*r^2*7

OpenStudy (danjs):

find r= radius and double it to get diameter

OpenStudy (crashonce):

the formula for the volume of a cylinder is stated by @DanJS : V=pir^2 h

OpenStudy (dtan5457):

175 = pi*r^2*7 yea, so from here find what makes r^2 times 7, equal 175, can you do that?

OpenStudy (dtan5457):

yep^ the radius is 5, which 5^2=25 times 7=175 diameter is twice the radius, so it's 10

OpenStudy (anonymous):

thank you all:)

OpenStudy (danjs):

i got the diameter is 5.64

OpenStudy (danjs):

\[\frac{ 10 }{ \sqrt{\pi} } \approx 5.64\]

OpenStudy (danjs):

oh 175 pi is the volume, not just 175

OpenStudy (danjs):

yeah then diameter is 10

OpenStudy (danjs):

\[175 \pi = \pi * r^2 * 7\] \[175 = r^2 *7\] \[r = 5\] d=2r = 10